OpenStudy (anonymous):

Solve by ELIMATION. 2x-3y=61 2x+y=-7

jimthompson5910 (jim_thompson5910):

Multiply everything in equation 2 by -1 to go from 2x+y=-7 to -2x-y=7

jimthompson5910 (jim_thompson5910):

So you'll have this new system 2x-3y=61 -2x-y=7 Now add the equations to get 2x-3y=61 -2x-y=7 ---------- 0x - 4y = 68 so this means -4y = 68 y = ???

OpenStudy (anonymous):

y=-17

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now use this to find x

jimthompson5910 (jim_thompson5910):

pick any equation with x and y, plug in y = -17 and solve for x 2x-3y=61 2x-3(-17)=61 2x + 51 = 61 ... ... ... x = ???

OpenStudy (anonymous):

x=-5

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

I meant positive 5. What about this problem: 2x+18y=-9 & 4x+18y= -27

jimthompson5910 (jim_thompson5910):

good, x = 5, y = -17

jimthompson5910 (jim_thompson5910):

notice how we have 18y in each equation

jimthompson5910 (jim_thompson5910):

if we add them we get 18y + 18y = 36y which isn't 0y but if we change the sign of either 18y, we will be able to add them so we get 0 since 18y + (-18y) = 0y = 0

jimthompson5910 (jim_thompson5910):

see how this works?

OpenStudy (anonymous):

So would we multiply it by -1?

jimthompson5910 (jim_thompson5910):

yes, either equation (both sides of it)

OpenStudy (anonymous):

Would x equal 9?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

What would it be then?

jimthompson5910 (jim_thompson5910):

if you multiplied both sides of equation two by -1, then you would have 2x+18y=-9 -4x-18y= 27

jimthompson5910 (jim_thompson5910):

adding the equations gives you 2x+18y=-9 -4x-18y= 27 ------------ -2x + 0y = 18 so, -2x = 18 ----> x = ??

OpenStudy (anonymous):

Okay I have that.

jimthompson5910 (jim_thompson5910):

but x is not 9

jimthompson5910 (jim_thompson5910):

since -2x = 18 -2(9) = 18 -18 = 18 is not true

jimthompson5910 (jim_thompson5910):

so what is x instead?

OpenStudy (anonymous):

I don't get it..

jimthompson5910 (jim_thompson5910):

-2x = 18 divide both sides by -2 to isolate x x = 18/(-2) x = -9

jimthompson5910 (jim_thompson5910):

so you were close, just off by a sign

jimthompson5910 (jim_thompson5910):

now use this to find y

OpenStudy (anonymous):

I even had that haha I just didn't see the sign

jimthompson5910 (jim_thompson5910):

oh lol, that's good

jimthompson5910 (jim_thompson5910):

anyways, what do you get for y?

OpenStudy (anonymous):

y=1/2

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Okay next one, 3x-10y=-25 & 4x+40y=20

jimthompson5910 (jim_thompson5910):

are any of the x or y coefficients equal? or equal and opposite in sign?

OpenStudy (anonymous):

Negative 10 and positive 40 are opposites in signs

jimthompson5910 (jim_thompson5910):

how can you change the -10y to -40y ?

jimthompson5910 (jim_thompson5910):

so you can add -40y and +40y to get 0y

OpenStudy (anonymous):

Multiply -10 by positive 4

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you have to do that to both sides of everything you change in the equation that has -10y

jimthompson5910 (jim_thompson5910):

so multiply both sides of 3x-10y=-25 by 4 to get 3x-10y=-25 4(3x-10y)=4(-25) 12x - 40y = -100

jimthompson5910 (jim_thompson5910):

So 3x-10y=-25 4x+40y=20 turns into 12x - 40y = -100 4x+40y=20

OpenStudy (anonymous):

Okay would you add or subtract?

jimthompson5910 (jim_thompson5910):

add because -40y + 40y = 0y

OpenStudy (anonymous):

okay would x=-5?

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

And y=1

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Okay two more left. x-8y-18 & -16x+16y=-8

OpenStudy (anonymous):

Would you multiply the first one by -2

jimthompson5910 (jim_thompson5910):

x-8y-18 looks like a typo

OpenStudy (anonymous):

I meant =18

jimthompson5910 (jim_thompson5910):

ok so x-8y=18

jimthompson5910 (jim_thompson5910):

if you multiply both sides by 2, you get 2x-16y = 36

jimthompson5910 (jim_thompson5910):

x-8y=18 -16x+16y=-8 turns into 2x-16y = 36 -16x+16y=-8

OpenStudy (anonymous):

I thought it would -2x-16y=36

jimthompson5910 (jim_thompson5910):

why -2x ?

OpenStudy (anonymous):

And wouldnt the 36 be negative too

jimthompson5910 (jim_thompson5910):

18 is positive, so multiplying +2 by +18 gives you +36

OpenStudy (anonymous):

I thought we were multiplyinh it by -2 sorry.

jimthompson5910 (jim_thompson5910):

oh i see, no we're using +2

jimthompson5910 (jim_thompson5910):

if we used -2, then -8y would turn into 16y

jimthompson5910 (jim_thompson5910):

we want it to be -16y so the -16y and +16y add to 0y

jimthompson5910 (jim_thompson5910):

the terms have to have equal coefficients but opposite signs so they add to 0

OpenStudy (anonymous):

It would be x=-2

jimthompson5910 (jim_thompson5910):

x = -2 is correct

OpenStudy (anonymous):

Any y= -2 1/2

jimthompson5910 (jim_thompson5910):

or y = -5/2, nice work

OpenStudy (anonymous):

Would 88x-5y=39 & -8x+3y=-1 be 0,7 4/5

jimthompson5910 (jim_thompson5910):

no assuming there are no typos in the original equations

OpenStudy (anonymous):

Will you walk me through it?

jimthompson5910 (jim_thompson5910):

we have 88x and -8x what can we do to -8x to get it to -88x?

OpenStudy (anonymous):

Multiply it by 11

jimthompson5910 (jim_thompson5910):

so you have to multiply both sides of equation 2 by 11

jimthompson5910 (jim_thompson5910):

that means 88x-5y=39 -8x+3y=-1 turns into 88x-5y=39 -88x+33y=-11 Add the equations to get: 88x-5y=39 -88x+33y=-11 ------------ 0x + 28y = 28 which means 28y = 28 so what is the value of y?

OpenStudy (anonymous):

y=1

jimthompson5910 (jim_thompson5910):

use that to find x

OpenStudy (anonymous):

x=1/2

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

Thank you so much for your help! I really appreciate it!

jimthompson5910 (jim_thompson5910):

sure thing